OpenStudy (anonymous):

you are smart if you solved this: A CURVE HAS THE PARAMETRIC EQUATIONS OF; X=t-3/t AND Y=t^2+4/t DETERMINE; i ). dy/dx .., ii ). d^2y/dx^2 how much time you'll take to do it.. please right down the time you took when you solve it.

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=\frac{ \frac{ dy }{ dt } }{ \frac{ dx }{ dt } }\] dx/dt= 1+3/t^2 dy/dt= 2t-4/t^2 So \[\frac{ dy }{ dx }=\frac{ {2t -\frac{ 4 }{ t^2 }} }{{1+\frac{ 3 }{ t^2 }} }\]

OpenStudy (anonymous):

so, you still remember how to do that, congratulation for doing it.

OpenStudy (anonymous):

I did double check online before doing it

OpenStudy (anonymous):

but u did it anyway